Find a function whose graph is the line segment joining the points (-5, -3) and (10, 5).
f(x)=________

Answer by Alan3354(69443)

|+x +y 1|
|-5 -3 1| = 0
|10 +5 1|
x*(-3-5) - y*(-5-10) + 1*(-25+30) = 0
-8x + 15y + 5 = 0
8x - 15y = 5
or, y = (8/15)x - 1/3
f(x) = (8/15)x - 1/3